# An Agnostic Biosignature Based on Modeling Panspermia and Terraformation

### v1

@@ Abstract
A fundamental goal of astrobiology is to detect life outside of Earth. This proves to be an exceptional challenge outside of our solar system, where strong assumptions must be made about how life would manifest and interact with its planet. Such assumptions are required because of the lack of a consensus theory of living systems, or an understanding of the possible extent of planetary dynamics. Here we explore a model of life spreading between planetary systems via panspermia and terraformation. Our model shows that as life propagates across the galaxy, correlations emerge between planetary characteristics and location, and can function as a population-scale agnostic biosignature. This biosignature is agnostic because it is independent of strong assumptions about any particular instantiation of life or planetary characteristic—by focusing on a specific hypothesis of what life may do, rather than what life may be. By clustering planets based on their observed characteristics, and examining the spatial extent of these clusters, we demonstrate (and evaluate) a way to prioritize specific planets for further observation—based on their potential for containing life. We consider obstacles that must be overcome to practically implement our approach, including identifying specific ways in which better understanding astrophysical and planetary processes would improve our ability to detect life. Finally, we consider how this model leads us to think in novel ways about hierarchies of life and planetary scale replication.
@@ 1 Introduction
The feasibility of interstellar lithopanspermia (non-intelligent exchange of material via rocks) has been discussed at length in other work (see e.g. Carroll-Nellenback et al. (2019); Grimaldi et al. (2021); Gobat (2021); Totani (2023) and references therein). Crucially, it appears plausible, although likelihoods and rates vary considerably depending on assumptions on timing, amount of material ejected, organismal hardiness, and capture rates, among other features. Arguments have likewise been made for the feasibility of directed interstellar panspermia by intelligent life (Wright et al., 2014). Ultimately, our postulates of panspermia and terraformation are merely well understood hallmarks of life (proliferation via replication, and adaptation with bi-directional environmental feedback), escalated to the planetary scale, and executed on an interstellar scale.
@@ 2.2.2 Mantel Test
The Mantel test is a measure of correlation between two distance matrices. The resulting value is called the Mantel coefficient, and is reported alongside a p-value calculated from an approximate permutation test (indicating the proportion of randomly permuted distance matrices which have correlations greater or equal to the correlation between the non-permuted distance matrices). Simply put, the p-value indicates how unlikely this correlation is compared to random permutations of the data. We wrote the Mantel test in Julia, with the algorithm and code adapted from Python’s scikit-bio (The scikit-bio development team, 2020; Rideout et al., 2023). Here we used the Mantel test to measure the Pearson correlation between a distance matrix of all planet positions, and a distance matrix of all planet compositions. For each Mantel coefficient calculation, the p-value was generated using permutations with a 2-sided alternative hypothesis. This is approximately equal to confidence. Because the p-value quantifies how anomalous an observed composition/position association is, given the assumptions of the model that correlations should only occur from panspermia and terraformation, we treat the p-value as one measure of confidence in the space containing a biosignature. To get an idea of how sensitive the Mantel coefficient corresponding to a detection is, we plotted how it varies based on number of planets observed (Fig. A3). We find that the sensitivity of the Mantel coefficient to number of planets observed decreases exponentially, and 1000 planets seems like a reasonable choice to reflect the balance of the challenge of realistically observing planets, with the need for those planets to exhibit potentially small correlations in composition-position space. The exact shape of this plot will vary by model parameters, but is especially dependent on the distribution of planet compositions and positions (e.g., planets being evenly distributed in composition or position space, vs. extremely heterogeneous).
@@ 2.2.4 Selecting clusters
We created a threshold for the spatial spread of the planets in the clusters by taking the average of the interquartile range (IQR, where the middle 50% of planets fall) of all planets across each of the dimensions, , within the spatial extent of a cube with sides of length 1/2 the size of the model space (i.e., with a volume = the model space). For the simulation analyzed in this paper, . Thus, any cluster with meets our threshold for being spatially localized666Though the extent we chose for an IQR threshold is arbitrary, it reflects the relative size of our model space and the presumption that looking for life that is still relatively spatially localized is of greater interest than looking for life which has already spread over the galaxy (the latter implying there might be other easier ways to look for life). This choice, like others, could be modified depending on other assumptions or objectives. .

### v2

@@ Abstract
A fundamental goal of astrobiology is to detect life outside of Earth. This proves to be an exceptional challenge outside of our solar system, where strong assumptions must be made about how life would manifest and interact with its planet. Such assumptions are required because of the lack of a consensus theory of living systems, or an understanding of the possible extent of planetary dynamics. Here we explore a model of life spreading between planetary systems via panspermia and terraformation. Our model shows that as life propagates across the galaxy, correlations emerge between planetary characteristics and location, and can function as a population-scale agnostic biosignature. This biosignature is agnostic because it is independent of strong assumptions about any particular instantiation of life or planetary characteristic—by focusing on a specific hypothesis of what life may do, rather than what life may be. By clustering planets based on their observed characteristics, and examining the spatial extent of these clusters, we demonstrate (and evaluate) a way to prioritize specific planets for further observation—based on their potential for containing life. We consider obstacles that must be overcome to practically implement our approach, including identifying specific ways in which better understanding astrophysical and planetary processes would improve our ability to detect life. Finally, we consider how this model leads us to think in novel ways about scales of life and planetary replication.
@@ 1 Introduction
The feasibility of interstellar lithopanspermia (non-intelligent exchange of material via rocks) has been discussed at length in other work (see e.g. Carroll-Nellenback et al. (2019); Grimaldi et al. (2021); Gobat (2021); Totani (2023) and references therein). Crucially, it appears plausible, although likelihoods and rates vary considerably depending on assumptions on timing, amount of material ejected, organismal hardiness, and capture rates, among other features. Arguments have likewise been made for the feasibility of directed interstellar panspermia by intelligent life (Wright et al., 2014; Newman & Sagan, 1981)Ultimately, our postulates of panspermia and terraformation are merely well understood hallmarks of life (proliferation via replication, and adaptation with bi-directional environmental feedback), escalated to the planetary scale, and executed on an interstellar scale.
@@ 2.2.2 Mantel Test
The Mantel test666While originally developed for biostatistics, and often applied to ecological data, we use the Mantel test here because it is a straightforward and established way to measure the correlation between a distance matrix representing space, and one representing an arbitrary attribute. The Mantel test has both critics and advocates (Legendre et al., 2015; Somers & Jackson, 2022) , which we mention simply to make clear that the approach we describe is not wedded to using the Mantel test—we would expect similar results when using other distance-based tests. is a measure of correlation between two distance matrices. The resulting value is called the Mantel coefficient, and is reported alongside a p-value calculated from an approximate permutation test (indicating the proportion of randomly permuted distance matrices which have correlations greater or equal to the correlation between the non-permuted distance matrices). Simply put, the p-value indicates how unlikely this correlation is compared to random permutations of the data. We wrote the Mantel test in Julia, with the algorithm and code adapted from Python’s scikit-bio (The scikit-bio development team, 2020; Rideout et al., 2023). Here we used the Mantel test to measure the Pearson correlation between a distance matrix of all planet positions, and a distance matrix of all planet compositions. For each Mantel coefficient calculation, the p-value was generated using permutations with a 2-sided alternative hypothesis. This is approximately equal to confidence. Because the p-value quantifies how anomalous an observed composition/position association is, given the assumptions of the model that correlations should only occur from panspermia and terraformation, we treat the p-value as one measure of confidence in the space containing a biosignature. To get an idea of how sensitive the Mantel coefficient corresponding to a detection is, we plotted how it varies based on number of planets observed (Fig. A3). We find that the sensitivity of the Mantel coefficient to number of planets observed decreases exponentially, and 1000 planets seems like a reasonable choice to reflect the balance of the challenge of realistically observing planets, with the need for those planets to exhibit potentially small correlations in composition-position space. The exact shape of this plot will vary by model parameters, but is especially dependent on the distribution of planet compositions and positions (e.g., planets being evenly distributed in composition or position space, vs. extremely heterogeneous).
@@ 2.2.4 Selecting clusters
We created a threshold for the spatial spread of the planets in the clusters by taking the average of the interquartile range (IQR, where the middle 50% of planets fall) of all planets across each of the dimensions, , within the spatial extent of a cube with sides of length 1/2 the size of the model space (i.e., with a volume = the model space). For the simulation analyzed in this paper, . Thus, any cluster with meets our threshold for being spatially localized777Though the extent we chose for an IQR threshold is arbitrary, it reflects the relative size of our model space and the presumption that looking for life that is still relatively spatially localized is of greater interest than looking for life which has already spread over the galaxy (the latter implying there might be other easier ways to look for life). This choice, like others, could be modified depending on other assumptions or objectives. .
@@ 4 Summary and Discussion
Applying our approach to observed astronomical data will also be affected by relative stellar velocities. Stars move relative to each other fast enough that, for example, the Sun’s current nearest neighbor, Alpha Centauri, will be on the opposite side of the galaxy on time scales on the order of 100 My. Naturally this is a problem if we wish to detect potential correlations on the timescale of the evolution of planets. In the worst case scenario, the approach we provide would be applicable for stars in similar phase space (Lin & Loeb, 2015), or for life which spreads much faster than relative stellar velocities—potentially limiting its applicability to life spread via directed panspermia, or via lithopanspermia on exceptionally fast moving interstellar objects.